OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (yttrium):

y = mx+b, where m = slope and b = y intercept

OpenStudy (anonymous):

these are the answers i can chose from A:slope = 4/5 and y-intercept = −9/5 B:slope = −4/5 and y-intercept = 9/5 C:slope = −4/5 and y-intercept = −9/5 D:slope = 4/5 and y-intercept = 9/5

OpenStudy (yttrium):

So, what will be your answer then?

OpenStudy (texaschic101):

do like Yttrium said...put it in y = mx + b form where m is the slope and y is the y intercept.

OpenStudy (anonymous):

B right Because of the -9

OpenStudy (texaschic101):

oops..b is the y intercept

OpenStudy (yttrium):

@rezo, can you show us first the form y = mx + b of our given equation?

OpenStudy (texaschic101):

B is not correct

OpenStudy (anonymous):

y=4x+5 Right

OpenStudy (texaschic101):

4x + 5y = -9 subtract 4x from both sides and what do you get ?

OpenStudy (yttrium):

No. It's actually y = -4x/5 - 9/5

OpenStudy (anonymous):

So the slope is -4/5 and the y interseppt is -9/5

OpenStudy (texaschic101):

correct

OpenStudy (anonymous):

That is Awsome thanks

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3?

OpenStudy (yttrium):

It's the converse. :D

OpenStudy (yttrium):

y - y1 = m (x-x1) substitute the values and you can find the equation. :)